jean is twice as old as bob. the sum of thier ages 3 years ago was 45 years. what are thier ages now?

Found 2 solutions by Superownage, chitra:
Answer by Superownage(7)

Hello, this one can be a bit of a problem. Add three years for both of the people, so 51 is the new total. Now jean is 2x (2 times as old as bob) and bob is x, there are three parts of this so 51/3 = bob is 17 and jean is 34

Answer by chitra(359)

Let bobs age be "x" yrs.

So, therefore Jeans age = 2x

Given that sum of their ages 3 yrs ago was 45 yrs

(x + 2x) - 3 = 45

3x - 3 = 45

3(x - 1) = 45

(x - 1) = 45/3

x - 1 = 15

x = 15 + 1

x = 16 yrs

This is Bobs age

So therefore Jeans age is:

2(x) = 2(16)

= 32 yrs

Jeans age is 32 yrs.

these are their current ages.


Regs